OpenStudy (lgbasallote):

Locate the fallacy: Let a= b Then a^2 = ab a^2 + a^2 = a^2 + ab 2a^2 = a^2 + ab 2a^2 - 2ab = a^2 + ab - 2ab 2a^2 - 2ab = a^2 - ab 2(a^2 - ab) = a^2 - ab --------- ------- a^2 - ab a^2 - ab 2 = 1

OpenStudy (lgbasallote):

This is a classic easy one :DDD

OpenStudy (anonymous):

dividing by a^2-ab, if a=b, then a^2-ab=ab-ab, division by zero

OpenStudy (unklerhaukus):

the division is wrong, you are dividing by zero

OpenStudy (lgbasallote):

hahahaha you guyzzz are gooood =)))))

OpenStudy (turingtest):

http://9gag.com/gag/689049

OpenStudy (anonymous):

saw this yesterday in general math errors..:P:P:P

OpenStudy (lgbasallote):

congrats on the modship btw @TuringTest :DDD

OpenStudy (turingtest):

Thanks @Igbasallote :D